OpenStudy (anonymous):

In an isosceles triangle, the length of each leg exceeds twice the base by 12cm, and the perimeter is 105cm. How long is each side?

OpenStudy (ranga):

Assume each side is x and the base is b. x = 2b + 12 ---- (1) Perimeter is: x + x + b = 105 2x + b = 105 ------ (2) Two equations, two unknowns. Solve for b.

OpenStudy (anonymous):

need more help understanding.

OpenStudy (ranga):

Let me know which part is not clear. The first equation? The second equation? Or how to solve 2 equations with 2 unknowns?

OpenStudy (anonymous):

Like what do I put in for b?

OpenStudy (ranga):

They are asking for x. To find x you have to solve the two equations: x = 2b + 12 or x - 2b = 12 2x + b = 105 Multiply second equation by 2: 4x + 2b = 210. Add it to the first equation x - 2b = 12 add 5x = 222 x = 222/5 = 44.4 cm.

OpenStudy (anonymous):

Thank u so much!